# The Impacts of the Gender Imbalance on the Marriage Market:

### v3

@@ 1 Introduction
War causes enormous losses of people. These losses are generally concentrated on men, who are drafted into battle, rather than women, leading to a substantial reduction in the sex ratio in an economy (i.e., the ratio of men to women in the population). Theoretically, relative male scarcity improves the bargaining position of men in the marriage market and thus their intra-household allocations (Becker 1973, 1974, 1991; Chiappori et al. 2002). A growing body of empirical research has validated this prediction, providing solid evidence that the wartime losses of men affect not only marriage market outcomes but also birth outcomes both inside and outside of marriage (Abramitzky et al. 2011; Bethmann and Kvasnicka 2012; Brainerd 2017). [^1] See also Acemoglu and Autor (2004) for the impacts of war-induced changes in the female labor supply on earnings inequalities. Another strand of the literature exploits different semi-experimental approaches. For example, Angrist (2002), Lafortune (2013), and Porter (2015) use exogenous changes in the sex ratio due to the inflow of immigrants and famine to analyze the impacts on pre-marital investment, marriage and labor markets, consumption behavior, and the health status of offspring. Charles and Luoh (2010) analyze the impacts of male incarceration rates on women in the marriage market. Wei and Zhang (2011) and Edlund et al. (2013) employ the culture-induced gender imbalance to investigate the impacts on saving behavior and crime rates in China, respectively.
This study examines the impacts of relative male scarcity caused by World War II (hereafter WWII) on marriage market outcomes in postwar Japan. Japan’s wartime loss of men led to an unprecedented decline in the sex ratio, similar to that in post-World War I (hereafter WWI) France (Abramitzky et al. 2011), and post-World War II Bavaria, Germany (Bethmann and Kvasnicka 2012), and Russia (Brainerd 2017). We digitized census-based historical statistics and applied an instrumental variable (IV) technique that utilized exogenous variations in military mortality. We found that women facing relative male scarcity are less likely to marry. By contrast, men were able to marry even under extreme male losses. This clear sex difference suggests that men had a stronger bargaining position in the marriage market at that time. In addition, our results show that the estimated magnitude of women’s marriages decreased moderately between 1950 and 1955 when younger cohorts with a natural gender balance entered the marriage markets. We also find that widowed women facing relative male scarcity are less likely to remarry, and this tendency becomes clearer after the reinstatement of the military pension for widows in 1953.
@@ 2.1 Gender Imbalance
As many countries, Japan lost a number of men during WWII: $1,864,710$ military personnel died or were missing in action during the war (Nakamura and Miyazaki 1995, p. 289). [^3] This figure includes victims who died or were missing between 1942 and 1948. Deaths due to execution and diseases contracted on the frontline as well as those during the early stage of the Sino-Japanese war are not included (Nakamura and Miyazaki 1995, p. 289). Thus, the overall death toll was greater than the figure reported herein. A survey conducted in May 1948 shows that 323,495 homefront people died or were missing mainly due to bombing (Nakamura and Miyazaki 1995, p. 277). To understand the magnitude and persistence of the wartime losses of men, we first digitize the number of men and women using population censuses conducted after the war. Figure 1 illustrates the national average sex ratios by age in 1947, 1950, 1955, and 1960. The distributions of these sex ratios are considerably different from those in the prewar period (Online Appendix Figure B.2 illustrates the sex ratios in 1930 and 1935). This means that the clear reductions in the sex ratios in Figure 1 were caused by the wartime losses of men. We can highlight a few important features in those figures.
@@ 2.2 Marriages, Divorces, and Widowhoods
Figure 2a illustrates the marriage and divorce rates in the census years, indicating a clear hump in 1950.555This does not simply represent the decline in the population but rather the rises in the number of marriages and divorces. The numbers of marriages and divorces in 1935 are 556,730 and 48,528, respectively, whereas those are 715,081 and 83,680 in 1950 and 714,861 and 75,267 in 1955 (Statistics Bureau of the Cabinet 1936; Division of Health and Welfare Statistics, Welfare Minister’s Secretraiat 1953, 1957). A large number of people who could not marry during the war started to marry thereafter, creating a clear marriage boom in the early 1950s (Yuzawa 1977). Correspondingly, the average age at first marriage also decreased in 1950, as shown in Figure 2b. [^6] Some people in their early 20s during the war married in the wartime period under the pronatalist policy (called kakekomi kon, meaning hasty marriages), which might have led to the declines in average age at first marriage (Toshitani 1984). However, these marriages did not lead to a clear increment in the national average marriage rate (Statistics Bureau of the Cabinet 1939a, 1936).
@@
Overall, these figures suggest that relative male scarcity due to the war might strengthen men’s bargaining position in the marriage market. Indeed, an article in a popular magazine titled Ie-no-Hikari (light of a house) at that time claimed: “You (a widowed female) are not most likely to get remarried because there are a large number of women and a small number of men in the marriage market” (Kawaguchi 2003, pp. 124–125). In our empirical analysis, we consider information on the marriage status (single, married, divorced, and widowed) to comprehensively investigate the impact of men’s wartime losses on the marriage market.
@@ 3 Conceptual Framework
Alternative theoretical representation of marriage offers materially similar predictions about the effects of sex ratio changes. The unitary model proposed by Grossbard-Shechtman (1984) suggests that a decline in the sex ratio increases the labor force participation of females because of a decline in the shadow wage for household production. This means that females are less likely to marry under conditions of relative male scarcity. The collective model considering bargaining between wives and husbands suggests that a decline in the sex ratio decreases the bargaining position of the wife on intrahousehold resource allocation, given the better outside options for males (Chiappori et al. 2002). The framework of stable matching under the Gale-Shapley algorithm predicts that a decline in the sex ratio increases men’s welfare, but decreases that of women (Browning et al. 2014). Despite differences in framework, these theories suggest that the relative bargaining position of women could be lower in a society with relative male scarcity.
@@ 4 Data
Statistics on socioeconomic outcomes after the war are scarce, as in other countries. Fortunately, however, the 1950 and 1955 Population Censuses document prefecture-age-level information on marriage market outcomes and the number of people in Japan (Bureau of Statistics, Office of the Prime Minister 1951a, 1956a). We digitize these statistics to prepare data on the sex ratio and several measures of marriage market outcomes for 1950 and 1955.111111We digitize the data using 92 (46 editions for each census year) reports of the censuses in total. To conserve space, we display those as one citation (say, Bureau of Statistics, Office of the Prime Minister (1951a) and Bureau of Statistics, Office of the Prime Minister (1956a)) for each census year throughout this study. We cannot include Okinawa Prefecture in our analytical sample because the island remained under the exclusive control of the American military until May 1972. The Population Census is required by law and covers all the people living in areas under the administrative authority of the Japanese government on the survey date. This means that the census-based dataset is particularly useful for avoiding sample selection bias in statistical inference. In addition, the marriage status item was documented based on common-law marriage status instead of registration status. This means that cohabitation before marriage did not affect the statistics available in the census. [^12] See Sato (2015) for the statistics on cohabitations at that time.
@@
where MALE and FEMALE are the number of men and women, respectively. We use an age range spanning between 2 years younger to 10 years older for the following reasons. First, Japanese women were used to getting married to men aged 5–10 years older than themselves (Shiosawa and Shimada 1975, p. 11). Second, women occasionally married men who were a few years younger. One example is that of levirates, with Emori (1966) showing that they were prevalent in 1940s Japan. Therefore, it is practically reasonable to set the lower and upper limits of the thresholds as -2 and 10 for calculating the sex ratio. Online Appendix Figure B.1 illustrates the adjusted sex ratio by prefecture and age.
@@
We consider potential factors that might be correlated with the local sex ratios. First, we considered the possibility that regions with many military sites experience more male losses than other regions. Fortunately, the 1940 Population Census provides systematic statistics on the number of people working in the military factory (Bureau of Statistics Office of the Prime Minister 1961). We digitized the share of these workers among the total population measured in the same census to yield a proxy for the concentration of military sites in each prefecture at the initial stage. Similarly, we digitized homefront mortality due to wartime air attacks using an official report compiled by Nakamura and Miyazaki (1995). Both variables were expected to control for potential regional gender bias in wartime homefront casualties. Second, we digitized the mortality of the pioneer settlers of Manchuria using an official report named the List of Manchuria Pioneer Groups (Investigation Department for Unrepatriated Persons 1954). Because the number of pioneer settlers was considerably small relative to the homefront people, this should not disturb the local sex ratios. However, early settlers were more likely to be male than female, which is still preferable to be controlled for. In summary, our main results were not affected by controlling for these three variables. Despite this, to be conservative, we include all control variables in all the regressions. Online Appendices B.4B.5B.6 summarize the documents used to create the control variables.
@@ 5 Estimation Strategy
where MR is the military mortality rate, $\mathbf{x}$ is a vector of controls, $\nu$ is the age fixed effect, and $e$ is the random error term. Military mortality interacts with the indicator variable for affected cohorts (AC). The affected cohorts were the draft age cohorts during the war, which were exogenously determined by the draft order. [^14] Online Appendix C.1 provides a summary of the cohorts who were drafted under the wartime regime. Moreover, the drafts were conducted randomly during the war (Watanabe 2014a, b) and most of the changes in sex ratio around the war can be explained by the wartime casualties (Asai and Kanbayashi 2023). This indicates that military mortality is valid as an excluded exogenous variable and can reasonably predict the wartime losses of males in our system.
@@ 6 Results
Panel A of Table 2 presents the results for the 1950 sample, whereas Panel B shows those for the 1955 sample. Columns (1) to (4) list the estimates for the proportions of single, married, divorced, and widowed individuals. The first-stage $F$-statistics shown in both panels are reasonably high, not only for the 1950 but also for the 1955 sample. This supports evidence that the reduced-form equation satisfies the rank condition. Online Appendix C.2 summarizes the first-stage regression results. In addition, by employing control function regression, we confirmed that the exogeneity condition generally holds in our empirical setting. Section 7.3 presents these results.
@@
To summarize, relative male scarcity due to the war had economically meaningful impacts on women’s marriage. Women in regions with higher relative male scarcity, including widowhoods, were less likely to marry than women in other regions. In contrast, men suffered no clear impact from the war-induced gender imbalance, meaning that men were able to marry regardless of the degree of male scarcity. This result supports the evidence that men have a higher bargaining position in the marriage market after the war. We also found evidence that the estimated magnitude of the relative male scarcity in women’s marriages changed moderately in the 1955 sample. We interpret these findings in the Discussion section (Section 8). Before discussing the results, we assess the robustness of the main results in the next section.
@@ 8 Discussion
Unlike women, we found suggestive evidence that the males were less likely to be influenced by the war-induced gender imbalance. This means that men can marry and divorce, regardless of the distribution of local sex ratios. This seems reasonable because men are extremely scarce in the marriage market, and thus may easily marry in most regions (Section 2.2). [^18] As described in Section 2.3, while Japan achieved legal equality between men and women under the revised code of 1947, the traditional gender inequalities continued even after the war. This may also strengthen men’s bargaining positions in the local marriage market. The economic theory considers the relative change in the sex ratio, meaning that there are some cases in which women are scarce relative to men. However, in the case of war, all regions suffered absolute male losses. This feature in the shock on the local sex ratio may explain the sex differences in the results. Another interpretation may be related to the costs of divorce and marriage for men. In post-WW I France, the greater wartime losses of men decreased the proportion of divorced women and men, implying that women facing relative male scarcity were more likely to stay single and less likely to marry (Abramitzky et al. 2011, p. 136). Our results show that a similar bargaining mechanism existed in post-World War II Japan. However, as explained, marriage adjustments occurred outside of divorce in Japan. The case of Russia after World War II exemplifies this phenomenon, as men located in regions with greater male losses due to war were less likely to marry. Brainerd (2017) explains that the strong pronatalist Family Code of 1944, which led to the high cost of divorce as well as nearly costless non-marital sexual relations, decreased the probability of male marriage. In this light, Japanese men faced a similar situation because the revised code allowed divorce and division of properties at divorce (Section 2.3). This might have increased the costs of divorce and marriage for men, regardless of the degree of wartime male losses.
@@ 9 Conclusion
This study used male scarcity due to the casualties of WWII to analyze the impacts of the gender imbalance on marriage market outcomes. We found that women facing relative male scarcity were less likely to marry, whereas men were less likely to be influenced by the shock. While women’s overall situation in the marriage market did not improve throughout the 1950s, the reinterment of the military pension in 1953 might have reduced the incentive for remarriage among widowed women. Our evidence from post-World War II Japan has some limitations. While we investigate the middle-run (i.e., ten years after the war) effects of the gender imbalance on marriage market outcomes, we provide no evidence on the long-term impacts of the unbalanced sex ratio to deal with the potential influence of internal migration. The unavailability of systematic data on marriages by social status also made it difficult to analyze the impact of gender imbalance on assortative matching after the war. Despite these limitations, this study newly digitized a comprehensive census-based dataset with information on marriage market outcomes at two survey points and exploited plausibly exogenous variations in men’s wartime losses. This study provides suggestive evidence of the dynamic relationships between gender imbalance and changes in the marriage market as well as gender-based differences in its effects.

### v4

@@ 1 Introduction
War causes enormous losses of people. These losses are generally concentrated on men, who are drafted into battle, rather than women, leading to a substantial reduction in the sex ratio in an economy. Theoretically, relative male scarcity improves the bargaining position of men in the marriage market and their intra-household allocations (Becker 1973, 1974, 1991; Chiappori et al. 2002). A growing body of empirical research has validated this prediction, providing solid evidence that the wartime losses of men affect not only marriage market outcomes but also birth outcomes both inside and outside of marriage (Abramitzky et al. 2011; Bethmann and Kvasnicka 2012; Brainerd 2017). [^1] See also Acemoglu and Autor (2004) for the impacts of war-induced changes in the female labor supply on earnings inequalities. Another strand of the literature exploits different semi-experimental approaches. For example, Angrist (2002), Lafortune (2013), and Porter (2015) use exogenous changes in the sex ratio due to the inflow of immigrants and famine to analyze the impacts on pre-marital investment, marriage and labor markets, consumption behavior, and the health status of offspring. Charles and Luoh (2010) analyze the impacts of male incarceration rates on women in the marriage market. Wei and Zhang (2011) and Edlund et al. (2013) employ the culture-induced gender imbalance to investigate the impacts on saving behavior and crime rates in China, respectively.
This study examines the impacts of relative male scarcity caused by World War II (hereafter WWII) on marriage market outcomes in postwar Japan. Japan’s wartime loss of men led to an unprecedented decline in the sex ratio, similar to that in post-World War I (hereafter WWI) France, and post-World War II Bavaria, Germany, and Russia (Abramitzky et al. 2011; Bethmann and Kvasnicka 2012; Brainerd 2017). We digitized census-based historical statistics and applied an instrumental variable (IV) technique that utilized exogenous variations in military mortality. We found that women facing relative male scarcity are less likely to marry. By contrast, given the absolute male losses after the war, men could marry regardless of the relative male scarcity. This clear sex difference suggests that men had a stronger bargaining position in the marriage market at that time. In addition, our results show that the estimated magnitude of women’s marriages decreased moderately between 1950 and 1955 when younger cohorts with a natural gender balance entered the marriage markets. We also find that widowed women facing relative male scarcity are less likely to remarry, and this tendency becomes clearer after the reinstatement of the military pension for widows in 1953.
@@ 2.1 Gender Imbalance
As many countries, Japan lost a number of men during WWII: $1,864,710$ military personnel died or were missing in action during the war (Nakamura and Miyazaki 1995, p. 289). [^3] This figure includes victims who died or were missing between 1942 and 1948. Deaths due to execution and diseases contracted on the frontline as well as those during the early stage of the Sino-Japanese war are not included (Nakamura and Miyazaki 1995, p. 289). Thus, the overall death toll was greater than the figure reported herein. A survey conducted in May 1948 shows that 323,495 homefront people died or were missing mainly due to bombing (Nakamura and Miyazaki 1995, p. 277). To understand the magnitude and persistence of the wartime losses of men, we first digitize the number of men and women using population censuses conducted after the war. Figure 1 illustrates the national average sex ratios by age in 1947, 1950, 1955, and 1960. The distributions of these sex ratios are considerably different from those in the prewar period. [^4] Online Appendix Figure B.2 illustrates the sex ratios in 1930 and 1935. This means that the clear reductions in the sex ratios in Figure 1 were caused by the wartime losses of men. We can highlight a few important features in those figures.
@@ 2.2 Marriages, Divorces, and Widowhoods
Figure 2a illustrates the marriage and divorce rates in the census years, indicating a clear hump in 1950.666This does not simply represent the decline in the population but rather the rises in the number of marriages and divorces. The numbers of marriages and divorces in 1935 are 556,730 and 48,528, respectively, whereas those are 715,081 and 83,680 in 1950 and 714,861 and 75,267 in 1955 (Statistics Bureau of the Cabinet 1936; Division of Health and Welfare Statistics, Welfare Minister’s Secretraiat 1953, 1957). A large number of people who could not marry during the war started to marry thereafter, creating a clear marriage boom in the early 1950s (Yuzawa 1977). Correspondingly, the average age at first marriage also decreased in 1950, as shown in Figure 2b. [^7] Some people in their early 20s during the war married in the wartime period under the pronatalist policy (called kakekomi kon, meaning hasty marriages), which might have led to the declines in average age at first marriage (Toshitani 1984). However, these marriages did not lead to a clear increment in the national average marriage rate (Statistics Bureau of the Cabinet 1939a, 1936).
@@
Overall, these figures suggest that relative male scarcity due to the war might strengthen men’s bargaining position in the marriage market. Indeed, an article in a popular magazine titled Ie-no-Hikari (light of a house) claimed: “You (a widowed female) are not most likely to get remarried because there are a large number of women and a small number of men in the marriage market” (Kawaguchi 2003, pp. 124–125).
@@ 3 Conceptual Framework
Alternative theoretical representation of marriage offers materially similar predictions about the effects of sex ratio changes. The unitary model proposed by Grossbard-Shechtman (1984) suggests that a decline in the sex ratio increases the labor force participation of females because of a decline in the shadow wage for household production. This means that females are less likely to marry under conditions of relative male scarcity. The collective model considering bargaining between wives and husbands suggests that a decline in the sex ratio decreases the bargaining position of the wife on intrahousehold resource allocation, given the better outside options for males (Chiappori et al. 2002). The framework of stable matching under the Gale-Shapley algorithm predicts that a decline in the sex ratio increases men’s welfare, but decreases that of women (Browning et al. 2014). Despite differences in framework, these theories suggest that the relative bargaining position of women could be lower in a society with relative male scarcity. [^12] While this section focuses on the relationship between gender imbalance and bargaining position, Chiappori (2020) provides a recent and comprehensive review of broader frameworks for the marriage market.
@@ 4 Data
Statistics on socioeconomic outcomes after the war are scarce, as in other countries. Fortunately, however, the 1950 and 1955 Population Censuses document prefecture-age-level information on marriage market outcomes and the number of people in Japan (Bureau of Statistics, Office of the Prime Minister 1951a, 1956a). We digitize these statistics to prepare data on the sex ratio and several measures of marriage market outcomes for 1950 and 1955.131313We digitize the data using 92 (46 editions for each census year) reports of the censuses in total. To conserve space, we display those as one citation (say, Bureau of Statistics, Office of the Prime Minister (1951a) and Bureau of Statistics, Office of the Prime Minister (1956a)) for each census year throughout this study. We cannot include Okinawa Prefecture in our analytical sample because the island remained under the exclusive control of the American military until May 1972. The Population Census is required by law and covers all the people living in areas under the administrative authority of the Japanese government on the survey date. This means that the census-based dataset is particularly useful for avoiding sample selection bias in statistical inference. In addition, the marriage status item was documented based on common-law marriage status instead of registration status. This means that cohabitation before marriage did not affect the statistics available in the census. [^14] See Sato (2015) for the statistics on cohabitations at that time.
@@
where MALE and FEMALE are the number of men and women, respectively. We use an age range spanning between 2 years younger to 10 years older for the following reasons. First, Japanese women were used to getting married to men aged 5–10 years older than themselves (Shiosawa and Shimada 1975, p. 11). Second, women occasionally married men who were a few years younger. One example is that of levirates, with Emori (1966) showing that they were prevalent in 1940s. Therefore, it is practically reasonable to set the lower and upper limits of the thresholds as -2 and 10 for calculating the sex ratio. Online Appendix Figure B.1 illustrates the adjusted sex ratio by prefecture and age.
@@
We consider potential factors that might be correlated with the local sex ratios. First, we considered the possibility that regions with many military sites experience more male losses than other regions. Fortunately, the 1940 Population Census provides systematic statistics on the number of people working in the military factory (Bureau of Statistics Office of the Prime Minister 1961). We digitized the share of these workers among the total population measured in the same census to yield a proxy for the concentration of military sites in each prefecture at the initial stage. Similarly, we digitized homefront mortality due to wartime air attacks using an official report compiled by Nakamura and Miyazaki (1995). Both variables were expected to control for potential regional gender bias in wartime homefront casualties. The number of homefront deaths was considerably less than the number of military deaths overseas, implying that including these variables does not alter the results.
Second, we digitized the mortality of the pioneer settlers of Manchuria using an official report named the List of Manchuria Pioneer Groups (Investigation Department for Unrepatriated Persons 1954). Because the number of pioneer settlers was considerably small relative to the homefront people, this should not disturb the local sex ratios. However, early settlers were more likely to be male than female, which is still preferable to be controlled for. In summary, our main results were not affected by controlling for these three variables. Despite this, to be conservative, we include all control variables in all the regressions. Online Appendices B.4B.5B.6 summarize the documents used to create the control variables.
@@ 5 Estimation Strategy
where MR is the military mortality rate, $\mathbf{x}$ is a vector of controls, $\nu$ is the age fixed effect, and $e$ is the random error term. Military mortality interacts with the indicator variable for affected cohorts (AC). The affected cohorts were the draft age cohorts during the war, which were exogenously determined by the draft order. [^16] Online Appendix C.1 provides a summary of the cohorts who were drafted under the wartime regime. In addition, almost all wartime military deaths occurred in overseas battles, indicating that the occurrence of death involves sufficient randomness. Note, again, that Okinawa Prefecture, where land battles took place, is excluded from our analytical sample. See Committee to Edit on Air Raids in Japan (1981) for the casualties in Okinawa. Moreover, the drafts were conducted randomly during the war (Watanabe 2014a, b) and most of the changes in sex ratio around the war can be explained by the wartime casualties (Asai and Kanbayashi 2023). This indicates that military mortality is valid as an excluded exogenous variable and can reasonably predict the wartime losses of males in our system.
@@ 6 Results
Panel A of Table 2 presents the results for the 1950 sample, whereas Panel B shows those for the 1955 sample. Columns (1) to (4) list the estimates for the proportions of single, married, divorced, and widowed individuals. The first-stage $F$-statistics shown in both panels are reasonably high, not only for the 1950 but also for the 1955 sample. This supports evidence that the reduced-form equation satisfies the rank condition. Online Appendix C.2 summarizes the first-stage regression results. [^17] Online Appendix C.4 summarizes the results from the reduced-form regressions of equation 2. We confirmed that the IV estimates demonstrate much more stable results than those of the reduced-form regressions. In addition, by employing control function regression, we confirmed that the exogeneity condition generally holds in our empirical setting. Section 7.3 presents these results.
@@
To summarize, relative male scarcity due to the war had economically meaningful impacts on women’s marriage. Women in regions with higher relative male scarcity, including widowhoods, were less likely to marry than women in other regions. In contrast, men suffered no clear impact from the war-induced gender imbalance, indicating that men could marry regardless of the degree of male scarcity under the absolute male losses. This result supports the evidence that men have relatively high bargaining position in the marriage market after the war. We also found evidence that the estimated magnitude of the relative male scarcity in women’s marriages changed moderately in the 1955 sample. We interpret these findings in the Discussion section (Section 8).
@@ 8 Discussion
Unlike women, we found suggestive evidence that the males were less likely to be influenced by the war-induced gender imbalance. This indicates that, given the absolute male losses after the war, men could marry and divorce regardless of the distribution of local sex ratios. This seems reasonable because men are extremely scarce in the marriage market, and thus may easily marry in most regions (Section 2.2). [^21] As described in Section 2.3, while Japan achieved legal equality between men and women under the revised code of 1947, the traditional gender inequalities continued even after the war. This may also strengthen men’s bargaining positions in the local marriage market. The economic theory considers the relative change in the sex ratio, meaning that there are some cases in which women are scarce relative to men. However, in the case of war, all regions suffered absolute male losses. This feature in the shock on the local sex ratio may explain the sex differences in the results. Narrative evidence shows that men had an advantage in the marriage market, even in the prefecture where the decline in the sex ratio was marginal. [^22] For example, a newspaper article describes the marriage counselling situation in Toshima Ward of Tokyo prefecture as follows: “Since its opening, 30 men and women (37 men, 270 women) have come for consultation by 22nd March. As 80 people (26 men and 54 women) applied for counseling, it was decided to close the center for the time being and hold a group matchmaking session.” (Yomiuri Shimbun , 23rd March 1951). See also Yomiuri Shimbun (, 9th April 1949; 9th September 1950) and Chūō kōron sha (1949, 1st December 1949), a famous magazine for women, for similar articles. Another interpretation may be related to the costs of divorce and marriage for men. In post-WW I France, the greater wartime losses of men decreased the proportion of divorced women and men, implying that women facing relative male scarcity were more likely to stay single and less likely to marry (Abramitzky et al. 2011, p. 136). Our results show that a similar bargaining mechanism existed in post-World War II Japan. However, as explained, marriage adjustments occurred outside of divorce in Japan. The case of Russia after World War II exemplifies this phenomenon, as men located in regions with greater male losses due to war were less likely to marry. Brainerd (2017) explains that the strong pronatalist Family Code of 1944, which led to the high cost of divorce as well as nearly costless non-marital sexual relations, decreased the probability of male marriage. In this light, Japanese men faced a similar situation because the revised code allowed divorce and division of properties at divorce (Section 2.3). This might have increased the costs of divorce and marriage for men, regardless of the degree of wartime male losses.
@@ 9 Conclusion
This study used male scarcity due to the casualties of WWII to analyze the impacts of the gender imbalance on marriage market outcomes. We found that women facing relative male scarcity were less likely to marry, whereas men were less likely to be influenced by the shock. While women’s overall situation in the marriage market did not improve throughout the 1950s, the reinterment of the military pension in 1953 might have reduced the incentive for remarriage among widowed women. Our evidence from post-World War II Japan has some limitations. While we investigate the middle-run effects of the gender imbalance on marriage market outcomes, we provide no evidence on the long-term impacts of the unbalanced sex ratio to deal with the potential influence of internal migration. The unavailability of systematic data on marriages by social status also made it difficult to analyze the impact of gender imbalance on assortative matching after the war. Despite these limitations, this study newly digitized a comprehensive census-based dataset with information on marriage market outcomes at two survey points and exploited plausibly exogenous variations in men’s wartime losses. This study provides suggestive evidence of the dynamic relationships between gender imbalance and changes in the marriage market as well as gender-based differences in its effects.